# DIO: Refining Mutual Information and Causal Chain to Enhance Machine Abstract Reasoning Ability

### v4

@@ Abstract
Despite the outstanding performance of current deep learning models across various domains, their fundamental bottleneck in abstract reasoning remains unresolved. To address this challenge, the academic community has introduced Raven’s Progressive Matrices (RPM) problems as an authoritative benchmark for evaluating the abstract reasoning capabilities of deep learning algorithms, with a focus on core intelligence dimensions such as abstract reasoning, pattern recognition, and complex problem-solving. Therefore, this paper centers on solving RPM problems, aiming to contribute to enhancing the abstract reasoning abilities of machine intelligence. Firstly, this paper adopts a “causal chain modeling” perspective to systematically analyze the complete causal chain in RPM tasks: image $\rightarrow$ abstract attributes $\rightarrow$ progressive attribute patterns $\rightarrow$ pattern consistency $\rightarrow$ correct answer. Based on this analysis, the network architecture of the baseline model DIO is designed. However, experiments reveal that the optimization objective formulated for DIO, namely maximizing the variational lower bound of mutual information between the context and the correct option, fails to enable the model to genuinely acquire the predefined human reasoning logic. This is attributed to two main reasons: the tightness of the lower bound significantly impacts the effectiveness of mutual information maximization, and mutual information, as a statistical measure, does not capture the causal relationship between subjects and objects. To overcome these limitations, this paper progressively proposes three improvement methods:
@@
    Brando: By generating trainable hypothetical incorrect options, the Brando explicitly tightens the variational lower bound of mutual information.
@@
    WORLD: Abandoning the generative assumption, the WORLD constructs a Gaussian mixture model for the feature distribution of RPM images, providing an infinite and informative set of incorrect options through weighted sampling to further tighten the lower bound.
@@
    DIEGO: Introducing metadata supervision, DIEGO directly corrects semantic biases in the “abstract attributes $\rightarrow$ progressive patterns” link of the causal chain, aligning the model’s internal representations with human reasoning frameworks.
Experiments demonstrate that these three methods not only significantly enhance the accuracy of discriminative RPM tasks but also empower DIO to solve open-ended generative RPM problems. The findings of this paper offer causal-driven architectural design principles, learning objective refinement strategies, and novel insights into cross-modal generation for abstract reasoning research.
Index Terms:
Abstract Reasoning, Raven’s Progressive Matrices, Generative RPM, Variational Mutual Information, Causal Chain Modeling, Representation Learning.
@@ I Introduction
Deep learning, a cross-disciplinary technology that draws inspiration from the operational mechanisms of the human brain, has in recent years showcased formidable potential and transformative capabilities across numerous fields, with particularly outstanding accomplishments in generative tasks. Key technologies, which include Generative Adversarial Networks (GANs)[1], Variational Autoencoders (VAEs)[2], Diffusion Model[3], and the Transformer architecture that is based on the attention mechanism[4], empower models to generate high-quality and realistic images, texts, and videos from random noise or given conditions, thereby expanding the application scope of artificial intelligence in the realm of content creation.
In computer vision, deep learning, which makes use of deep convolutional neural networks and their variants, can efficiently extract and accurately classify image features, outperforming traditional methods in tasks such as object detection and image segmentation [5, 6, 7]. In the field of natural language processing, language models that are built on deep learning can automatically learn language grammar, semantics, and contextual associations, performing exceptionally well in tasks such as machine translation and text generation, and propelling the technology towards intelligence and automation [4, 8, 9]. In the multimodal cross-field of visual question answering systems, deep learning, which integrates visual and language information, can achieve joint understanding and reasoning, accurately answering natural language questions that are related to images [11].
Currently, research on Large Language Models (LLMs) is a focal point in the field of deep learning. These models, which possess a large number of parameters and are trained on massive data, demonstrate strong generalization ability in various natural language processing tasks, providing efficient and unified solutions for different tasks and facilitating the progress of artificial intelligence towards generalization and intelligence. However, the academic community has recently found that although LLMs have achieved remarkable results in language understanding and generation, they encounter obvious performance bottlenecks in abstract reasoning tasks[12, 13, 14, 15, 16]. Abstract reasoning ability, which is a vital component of human advanced intelligence and involves the understanding of abstract concepts, the inference of logical relationships, and the formulation of problem-solving strategies, serves as a key indicator for measuring the intelligence level of models. The shortcomings of LLMs not only expose the limitations of existing models in handling complex cognitive tasks but also underscore the significance of enhancing their abstract reasoning ability for achieving more powerful and general artificial intelligence.
Based on the above, this paper centers on the issue of improving the abstract reasoning ability of deep learning models. It explores effective model architecture design, training strategy optimization, and data augmentation methods, which aim to offer theoretical and practical support for breaking through the performance bottlenecks of existing models in abstract reasoning tasks and promoting the application and expansion of deep learning technology in a broader range of cognitive tasks.
@@ II Problem Description
In artificial intelligence, abstract reasoning ability stands as a core hallmark of intelligence and a critical prerequisite for achieving human-level or superhuman AI. Yet current AI technologies are constrained by heavy reliance on large-scale annotated data, performing poorly in abstract tasks requiring causal reasoning, such as medical diagnosis and financial analysis. While academic efforts have produced datasets like RAVEN [18] and PGM [19] for benchmarking reasoning capabilities, designing algorithms to handle symbolic logic, causality, and uncertainty remains challenging. Strengthening this ability demands algorithmic innovation integrated with interdisciplinary insights from cognitive science, neuroscience, and beyond, enabling breakthroughs through cross-domain collaboration.
@@ II-A Raven’s Progressive Matrices
Raven’s Progressive Matrices (RPM) [17] serves as an authoritative benchmark for evaluating the advanced cognitive capabilities of deep learning algorithms, focusing on core intelligence dimensions such as abstract reasoning, pattern recognition, and complex problem-solving. Its standardized testing framework demonstrates high correlation with human intelligence metrics, making it a reliable and stable tool for assessing the cognitive complexity of machine learning models. In current research, the RAVEN dataset [18] and PGM dataset [19] have emerged as the two most representative standardized test paradigms in the RPM task domain.
@@ II-B Related work
Image reasoning models such as CoPINet[21] and DCNet[22] primarily focus on learning differences and rules. Meanwhile, models like NCD[24], SAVIR-T[25], as well as neuro-symbolic systems (including PrAE, NVSA, ALANS[26, 27, 28]), have achieved promising results in both enhancing model interpretability and improving reasoning accuracy. RS-CNN and RS-TRAN[29] demonstrate exceptional capabilities in addressing problems related to RPM, while Triple-CFN[31] stands out among numerous models by implicitly extracting and indexing information about concepts and reasoning units, effectively enhancing reasoning accuracy.
@@ IV-D The Progressive Pattern Consistency Evaluation Module
As its name suggests, DIO’s progressive pattern consistency evaluation module is responsible for using a Transformer Encoder to compute the consistency among progressive pattern representations $P_{\alpha jkl}$ that share the same $j$-index and $l$-index. Specifically, the Transformer Encoder performs a dimension-preserving mapping on $\{P_{\alpha jkl}|k\in[1,9]\}$ to obtain $\{P^{\prime}_{\alpha jkl}|k\in[1,9]\}$. Repeating this for all $j\in[1,N]$ and $j\in[1,4]$ yields the complete representation set $\{P^{\prime}_{\alpha jkl}|j\in[1,N],k\in[1,9],l\in[1,4]\}$. Finally, each vector in this full set is mapped to a scalar via an MLP, producing the consistency score $s_{\alpha jkl}$. This process can be formally expressed as:
$\{P^{\prime}_{\alpha jkl}|k\in[1,9]\}$ $=\text{Trans-Encoder}(\{P_{\alpha jkl}|k\in[1,9]\})$ (3)
$s_{\alpha jkl}$ $=\text{MLP}(P^{\prime}_{\alpha jkl})$ (4)
The entire module can be represented as:
@@
$=\text{PCEM}(\{P_{\alpha jkl}|k\in[1,9]\}_{i})|_{j=1,l=1}^{N,4}$ (5)
@@
The softmax function then maps $\{S_{i}|i\in[8,16]\}$ to a probability distribution $\hat{P}(x|\{x_{i}\}_{i=1}^{8},\theta)$ parameterized by DIO:
@@
We introduce a loss function ${\ell}_{\text{DIO}}$ to refine the DIO-parameterized distribution $\hat{P}(x|\{x_{i}\}_{i=1}^{8},\theta)$, thereby enabling end-to-end training of all parameters across DIO’s modules:
${\ell}_{\text{DIO}}=-\mathbb{E}\left[\log\frac{\hat{P}(x=x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)}{\sum_{\hat{\alpha}=9,\hat{\alpha}\neq\alpha}^{16}\hat{P}(x=x_{\hat{\alpha}}|\{x_{i}\}_{i=1}^{8},\theta)}\right]$ (8)
where $\mathbb{E}[\cdot]$ denotes the expectation over the data distribution.
@@ V Deviation and Potential Correction Methods for Learning Objectives of End-to-End RPM-solving Model
This paper proposes that causal-chain-based architecture optimization can effectively enhance the performance of deep learning networks in solving RPM problems. However, further investigation reveals that this methodology still fails to fundamentally improve the logical reasoning capabilities of deep learning networks.
V-A The learning objective of the DIO
By abstracting the specific architecture of DIO (i.e., by merging equations (1)–(6)), we obtain a unified formulation that highlights its core inductive mechanism:
@@ V-A The learning objective of the DIO
By substituting Equation (7), DIO can instantiate the parameterized probability distribution $\hat{P}(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$:
@@
Notably, $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$, which denotes the mutual information between the context matrix $\{x_{i}|i\in[1,8]\}$ and the correct option $x_{\alpha}$, admits the following variational lower bound:
$I(x_{\alpha};\{x_{i}\}_{i=1}^{8})\geq\mathbb{E}\left[\log\frac{P(x_{\alpha}|\{x_{i}\}_{i=1}^{8})}{\frac{1}{K}\sum_{{\hat{\alpha}}\neq\alpha}P(x_{\hat{\alpha}}|\{x_{i}\}_{i=1}^{8})}\right]$ (11)
where $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8})$ represents the true probability that $x_{\alpha}$ can serve as the correct option for the given context matrix $\{x_{i}|i\in[1,8]\}$ and $K$ denotes the number of all potential incorrect options. If we consider $\hat{P}(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$ as a sufficient approximation of the true conditional distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8})$, then substituting Equation (10) into Equation (11) yields a mutual information lower bound parameterized by DIO:
@@
$I(x_{\alpha};\{x_{i}\}_{i=1}^{8})\geq\mathbb{E}\left[\log\frac{e^{\text{DIO}(\{x_{i},x_{\alpha}|i\in[1,8]\})}}{\frac{1}{K}\sum_{{\hat{\alpha}}\neq\alpha}e^{\text{DIO}(\{x_{i},x_{\hat{\alpha}}|i\in[1,8]\})}}\right],$ (12)
$\quad\quad\quad\quad\quad\quad\quad\text{where}\quad K=7$
It can be observed that the parameterized lower bound within Equation (12) aligns with DIO’s optimization objective, as expressed by the following equation:
@@
$\mathbb{E}\left[\log\frac{e^{\text{DIO}(\{x_{i},x_{\alpha}|i\in[1,8]\})}}{\frac{1}{K}\sum_{{\hat{\alpha}}\neq\alpha}e^{\text{DIO}(\{x_{i},x_{\hat{\alpha}}|i\in[1,8]\})}}\right]=-\ell_{\text{DIO}}+\log K$ (13)
Equation (13) reveals that DIO’s learning objective (i.e. maximizing $\ell_{\text{DIO}}$) is neither the human-like reasoning logic embedded in RPM problems nor even the suboptimal approach of directly maximizing mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$, but rather optimizing the variational lower bound of the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$.
Furthermore, based on the properties of logarithmic functions, the following inequality holds:
@@
$-{\ell}_{\text{DIO}}>\mathbb{E}\left[\log\frac{\hat{P}(x_{\alpha}|\{x_{i}\}_{i=1}^{8})}{\sum_{\hat{\alpha}=9}^{16}\hat{P}(x_{\hat{\alpha}}|\{x_{i}\}_{i=1}^{8})}\right]$ (14)
@@
The right-hand side of Inequality (14) is precisely the classical formulation for optimizing the probabilistic distribution $\hat{P}(x|\{x_{i}\}_{i=1}^{8},\theta)$ using the Cross-Entropy (CE) loss function. From this, it can be inferred that for those past end-to-end RPM-solving models that parameterize $\hat{P}(x|\{x_{i}\}_{i=1}^{8},\theta)$ and employ the CE loss function, their learning objectives essentially amount to optimizing a looser lower bound on the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$.
Equation (13) and Inequality (14) jointly uncover a pivotal fact: whether it’s the DIO model or previous end-to-end RPM-solving models, their learning objectives are essentially a lower bound of the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$. In this case, the looseness of the lower bound undoubtedly has a direct and significant impact on the performance of these models.
V-B Introducing Hypothetical Options to Derive a Tight Lower Bound for Mutual Information.
This paper contends that by utilizing DIO to parameterize a tighter variational lower bound and adjusting its optimization objective to maximize this new lower bound, we can achieve a more effective maximization of the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$, ultimately enhancing DIO’s performance.
By analyzing Formula (12), we deduce that increasing the number $K$ of incorrect options encountered by DIO can guide its learning objective toward a tighter variational lower bound. However, RPM problem instances only offer a limited number (seven) of incorrect options. To this end, this paper proposes to introduce additional hypothetical incorrect options for DIO. Specifically, we attempt to introduce a set of hypothetical options $\{x_{i}|i\in[17,\beta]\}$ into the training pipeline of DIO. Consequently, the conditional distribution $\hat{P}(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$ is reparameterized as:

### v7

@@ Abstract
Despite the outstanding performance of deep learning across various domains, their fundamental bottleneck in abstract reasoning remains unresolved. To tackle this challenge, the community has introduced Raven’s Progressive Matrices (RPM) as an authoritative benchmark for evaluating abstract reasoning, pattern recognition, and complex problem-solving capabilities. Centering on RPM, this paper aims to enhance the abstract reasoning ability of machine intelligence. This paper proposes a unified “causal-information” framework that systematically addresses three pivotal questions: (i) how to embed a human-like reasoning chain into network architecture; (ii) how to tighten the variational lower bound on mutual information, which limits existing models; and (iii) how to reduce semantic misalignment between learned representations and human-like causal chains. We deliver four incremental contributions:
@@
    DIO – a baseline that explicitly encodes the full causal chain “image $\rightarrow$ attributes $\rightarrow$ patterns $\rightarrow$ consistency $\rightarrow$ answer” and reveals that its learning objective, like that of previous end-to-end RPM solvers, is essentially a variational lower bound on the mutual information between contexts and solutions, thereby exposing how the tightness of this bound governs the performance of such models;
@@
    Brando – identifies that one key opportunity for tightening this bound is to supply additional “constructive incorrect options,” and designs a learnable mapper to provide such options, offering a novel strategy for tightening the variational lower bound on mutual information in reasoning models;
@@
    WORLD – models RPM image features with a Gaussian mixture, enabling unlimited, diverse, and target-sampled incorrect options that push DIO toward the theoretical bound and empower open-ended generative RPM solving;
  4. 4.
    DIEGO – rectifies semantic biases in the “attributes $\rightarrow$ patterns” link, aligning internal representations with human causal frameworks and boosting out-of-distribution (o.o.d.) generalization.
Experiments show that our complete model achieves significant performance gains on RAVEN, I-RAVEN, and PGM benchmarks, exhibits stronger generalization on several o.o.d. sub-tasks, and is capable of generating structurally diverse RPM solutions. This study not only achieves remarkable performance in abstract visual reasoning but also validates “causal-chain-modeling plus mutual-information-tightening” as a core design principle for AI.
@@ I Introduction
Deep learning, a cross-disciplinary technology inspired by the brain’s operational mechanisms, has demonstrated formidable potential across numerous fields and achieved outstanding accomplishments in generative tasks. Key technologies, including Generative Adversarial Networks (GANs)[1], Variational Autoencoders (VAEs)[2] and the attention-based Transformer architecture[3], enable models to generate high-quality images, texts and videos from random noise or given conditions, thereby expanding AI’s scope in content creation.
In CV, deep CNNs and their variants extract and classify features efficiently, surpassing traditional methods in detection and segmentation [4, 5, 6]. In NLP, deep language models learn grammar, semantics, and context, excelling in translation and generation [3, 7, 8]. In multimodal VQA, fused visual-linguistic representations enable joint reasoning to answer natural-language questions about images [10]. Large Language Models (LLMs), with massive parameters and data, now dominate deep learning by delivering unified, high-performance NLP solutions and driving AI toward generality. Yet recent studies [11, 12, 13] consistently show that LLMs encounter clear bottlenecks in abstract reasoning tasks that demand concept understanding, logical reasoning, and strategy formation, which are core hallmarks of human intelligence and key yardsticks for model capability. These shortcomings reveal current limits in tackling complex cognition and underscore the critical need to boost LLMs’ abstract reasoning capacity for stronger, more general artificial intelligence.
This paper focuses on enhancing deep learning’s abstract reasoning ability through effective model architecture design, training strategy optimization, and data augmentation methods that aim to provide both theory and practice to break current bottlenecks and extend deep learning to a broader perspective and cognitive tasks.
@@ II Problem Description
In AI, abstract reasoning is regarded as a core hallmark of intelligence and a key prerequisite for achieving human-level or superhuman AI. The research community has proposed Raven’s Progressive Matrices (RPM)[14] as an evaluation benchmark, thereby presenting the challenge of designing algorithms that simultaneously handle symbolic logic, causal modeling, and uncertainty.
@@ II-A Raven’s Progressive Matrices
Raven’s Progressive Matrices (RPM) [14] constitute an authoritative benchmark for assessing higher-order cognitive abilities in deep-learning systems, targeting core facets of intelligence such as abstract reasoning, pattern recognition and complex problem solving. Their standardized protocol correlates strongly with human intelligence metrics, offering a reliable gauge of model cognitive complexity. Today, the RAVEN [15] and PGM [16] datasets stand as the two most representative RPM testbeds.
@@ II-B Related work
Image reasoning models such as CoPINet[18] and DCNet[19] primarily focus on learning differences and rules. Meanwhile, models like NCD[21], SAVIR-T[22], as well as neuro-symbolic systems (including PrAE, NVSA, ALANS[23, 24, 25]), have achieved promising results in both enhancing model interpretability and improving reasoning accuracy. RS-CNN and RS-TRAN[26] demonstrate exceptional capabilities in addressing problems related to RPM.
@@ IV-D The Progressive Pattern Consistency Evaluation Module
As its name suggests, DIO’s progressive pattern consistency evaluation module is responsible for using a Transformer Encoder to compute the consistency among progressive-pattern representations $P_{\alpha jkl}$ that share the same $j$-index and $l$-index. Specifically, the Transformer Encoder performs a dimension-preserving mapping on $\{P_{\alpha jkl}|k\in[1,9]\}$ to obtain $\{P^{\prime}_{\alpha jkl}|k\in[1,9]\}$. Repeating this for all $j\in[1,N]$ and $l\in[1,4]$ yields the complete representation set $\{P^{\prime}_{\alpha jkl}|j\in[1,N],k\in[1,9],l\in[1,4]\}$. Finally, each vector in this full set is mapped to a scalar via an MLP, producing the consistency score $s_{\alpha jkl}$. This entire module can be represented as:
@@
$=\text{PCEM}(\{P_{\alpha jkl}|k\in[1,9]\})|_{j=1,l=1}^{N,4}$ (3)
@@
Subsequently, the softmax function maps the scores $\{S_{i}\mid i\in[8,16]\}$ derived by DIO to a probability distribution $\hat{P}(x\mid\{x_{i}\}_{i=1}^{8},\theta)$ parameterized by DIO:
@@
We introduce a loss function ${\ell}_{\text{DIO}}$ to refine this parameterized distribution $\hat{P}(x|\{x_{i}\}_{i=1}^{8},\theta)$, thereby enabling end-to-end training of all parameters across DIO’s modules:
${\ell}_{\text{DIO}}=\mathbb{E}\left[\log(\frac{\sum_{\begin{subarray}{c}{\hat{\alpha}}=9\\ {\hat{\alpha}}\neq\alpha\end{subarray}}^{16}\hat{P}(x_{\hat{\alpha}}|\{x_{i}\}_{i=1}^{8},\theta)}{\hat{P}(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)}+\delta)\right]-\log\delta$ (6)
where $\mathbb{E}[\cdot]$ denotes the expectation over the data distribution; $\delta$ is a small safeguard against numerical overflow, initialized at 1 and annealed to $1\times 10^{-5}$ as training progresses.
@@ V Deviation and Potential Correction Methods for Learning Objectives of End-to-End RPM-solving Model
In this section, we provide a theoretical analysis anchored in DIO’s optimization objective to assess the validity of the intuitively motivated methodology of causal-chain-based architecture optimization.
V-A The Learning Objective of DIO
By abstracting away the specific architecture of DIO (i.e., by collapsing Equations (1)–(4)), we obtain a unified formulation that highlights its core inductive mechanism:
@@ V-A The Learning Objective of DIO
Using Equation (5) and (7), we recast the parameterized distribution $\hat{P}(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$ into a form that explicitly embeds DIO’s inductive mechanism :
@@
Accordingly, the loss function $\ell_{\text{DIO}}$ can be rewritten as:
@@
$\ell_{\text{DIO}}=\mathbb{E}\left[\log\frac{\sum_{\begin{subarray}{c}{\hat{\alpha}}=9\\ {\hat{\alpha}}\neq\alpha\end{subarray}}^{16}e^{\text{DIO}(\{x_{i},x_{\hat{\alpha}}|i\in[1,8]\})}}{e^{\text{DIO}(\{x_{i},x_{\alpha}|i\in[1,8]\})}}\right]$ (9)
For the sake of discussion, we have temporarily removed the safeguard $\delta$.
Notably, $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$, which denotes the mutual information between the context matrix $\{x_{i}|i\in[1,8]\}$ and the correct option $x_{\alpha}$, admits the following variational lower bound [30]:
$I(x_{\alpha};\{x_{i}\}_{i=1}^{8})\geq-\mathbb{E}\left[\log\frac{\frac{1}{K}\sum_{{\hat{\alpha}}\neq\alpha}P(x_{\hat{\alpha}}|\{x_{i}\}_{i=1}^{8})}{P(x_{\alpha}|\{x_{i}\}_{i=1}^{8})}\right]$ (10)
Here, $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8})$ represents the true probability that $x_{\alpha}$ is the correct option given the context matrix $\{x_{i}|i\in[1,8]\}$, i.e., that $x_{\alpha}$ can serve as the solution. $K$ denotes the total number of potential incorrect options.
If we consider $\hat{P}(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$ as a reliable approximation of the true conditional distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8})$, then substituting Equation (8) into Equation (10) yields a mutual information lower bound parameterized by DIO:
@@
$I(x_{\alpha};\{x_{i}\}_{i=1}^{8})\geq-\mathbb{E}\left[\log\frac{\frac{1}{7}\sum_{\begin{subarray}{c}{\hat{\alpha}}=9\\ {\hat{\alpha}}\neq\alpha\end{subarray}}^{16}e^{\text{DIO}(\{x_{i},x_{\hat{\alpha}}|i\in[1,8]\})}}{e^{\text{DIO}(\{x_{i},x_{\alpha}|i\in[1,8]\})}}\right]$ (11)
It can be observed that the parameterized lower bound in Equation (11) matches DIO’s optimization objective in Equation (9):
@@
$-\mathbb{E}\left[\log\frac{\frac{1}{7}\sum_{\begin{subarray}{c}{\hat{\alpha}}=9\\ {\hat{\alpha}}\neq\alpha\end{subarray}}^{16}e^{\text{DIO}(\{x_{i},x_{\hat{\alpha}}|i\in[1,8]\})}}{e^{\text{DIO}(\{x_{i},x_{\alpha}|i\in[1,8]\})}}\right]=\log 7-\ell_{\text{DIO}}$ (12)
Equation (12) reveals that DIO’s learning objective (i.e., maximizing $\ell_{\text{DIO}}$) is neither human-like reasoning logic embedded in RPM problems nor even directly optimizing mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$, but rather maximizing a variational lower bound on that mutual information.
Furthermore, by the properties of logarithmic functions, the following inequality holds:
$-\log\hat{P}(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)>{\ell}_{\text{DIO}}$ (13)
@@
The left-hand side of Inequality (13) is precisely the expression for the loss function when DIO is optimized using the cross-entropy (CE) loss. Notably, inequality (13) abstracts away the specific feed-forward architecture of DIO, primarily retaining its inductive mechanism. This implies that the same inequality holds for any model parameterizing the distribution $\hat{P}(x|\{x_{i}\}_{i=1}^{8},\theta)$. Consequently, it can be inferred that for those prior end-to-end RPM-solving models trained with the CE loss function, their learning objectives amount to optimizing a looser lower bound on $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$.
Equation (12) and Inequality (13) jointly uncover a pivotal fact: whether training the DIO model using $\ell_{\text{DIO}}$ or previous end-to-end RPM solving models via the CE loss, their learning objectives are essentially maximizing lower bounds on the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$. As a result, regardless of how sophisticated the model architectures are, the tightness of these bounds directly and significantly determines the performance of such models. This observation further implies that in RPM problems the model architecture optimization is intrinsically limited; the crux is directing the model’s objective to maximize a tighter lower bound on the mutual information.
V-B Tightening the Lower Bound via Hypothetical Options
This paper contends that if DIO parameterizes a tighter variational lower bound and its learning objective explicitly targets this bound, we can more effectively maximize the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$, thereby further enhancing DIO’s performance. By analyzing Equation (10)–(12), we deduce that increasing the number $K$ of incorrect options presented to DIO can steer its learning objective toward a tighter variational lower bound. However, each RPM instance provides only seven such options. Given this, we propose supplying DIO with additional hypothetical incorrect options.
Specifically, we introduce a set of hypothetical options, denoted as $\{x_{i}|i\in[17,\beta]\}$, into the original RPM instance $\{x_{i}|i\in[1,16]\}$, allowing DIO to recast the parameterized distribution $\hat{P}(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$ from the form given in Equation (8) into the following: